OpenStudy (anonymous):

Solve the system of equations. 6x+15y=15000000 x+y=2000000

OpenStudy (arindameducationusc):

do you know to solve?

OpenStudy (anonymous):

I did and got x= 1666667 and y= 333333

OpenStudy (anonymous):

But I checked it but it was an approximate value

OpenStudy (anonymous):

To solve this first multiply second equation by 6 or 15 ok we multiply 2nd equation by 6 . Then subtract second equation from first, Then then u will get value of y , after that just put that value( of y) in any above equation u will get value of x...

OpenStudy (welshfella):

6x+15y=15000000 -6x - 6y= -12000000 adding 9y = 3,000,000 y = 333333.33

OpenStudy (welshfella):

x = 2000000 - 333333.33 = 1,666,666.67